OpenStudy (anonymous):

This helps us to simplify: square root of 96= square root (2x2) x (2x2) x2x3=4 square root 6 now simplify the square root of 72= square root ...... = square root ......

OpenStudy (anonymous):

\[\sqrt{72}=\sqrt{2^3 3^2}=\sqrt{2^2}\sqrt{2}\sqrt{3^2}=6\sqrt2\]